OpenStudy (anonymous):

If the line y=4x+3 is tangent to the curve y=x^2+c, then c is?

OpenStudy (anonymous):

differentiate the curve y, we'll get the slope of tangent which is 2x=4. so now we know that the tangent is at a point x=2 and y=4(2)+3=11. Substituting the value of this point (2,11) into the curve equation, 11=2^2+c thus c=7. :)

OpenStudy (anonymous):

how did you get 2x=4?

OpenStudy (anonymous):

the gradient of tangent line is 4 meanwhile the slope of tangent is 2x. they are the same :)

OpenStudy (anonymous):

ohhh ok...Thank you! :)

OpenStudy (anonymous):

welcome :)